# A Representation Learning Approach to Feature Drift Detection in Wireless Networks

### v1

@@ Acknowledgments
This work was funded in part from the Slovenian Research and Innovation Agency under the grant P2-0016 and in part from the European Union’s Horizon Europe Framework Program SNS-JU under grant agreement No 101096456 (NANCY). We also acknowledge that instead of using more traditional spell checking and language correction engines, we passed some of the manually written paragraphs in the introduction and related work, to AI asking for improvement of the flow and quality of the language.
@@
[Uncaptioned image] Blaž Bertalanič received his Ph.D. degree with highest distinction at the Faculty of Electrical engineering, University of Ljubljana. He is currently working as a researcher at Sensorlab, Jožef Stefan Institute. His main research interests are connected to advancement of machine learning and AI algorithms, especially in the context of time series analysis and smart infrastructures. Blaz is an IEEE member and with several leadership positions in the Slovenian chapter with over 15 IEEE publications including in IEEE Transactions on Neural Networks and Learning Systems.
[Uncaptioned image] George Floros received the Engineering Diploma, M.Sc., and Ph.D. degrees from the Department of Electrical and Computer Engineering, University of Thessaly, Volos, Greece, in 2013, 2015, and 2019, respectively. He is a Postdoctoral Researcher and an Academic Instructor with the Department of Electrical and Computer Engineering, University of Thessaly. His research interests lie primarily in the areas of mathematical modeling and simulation, model order reduction, as well as machine learning and data analysis.
@@
[Uncaptioned image] Carolina Fortuna is a research associate professor with the Jožef Stefan Institute where she leads Sensorlab. She was a Post-Doctoral Researcher with Ghent University, Ghent, Belgium. She was a Visiting researcher in Infolab with Stanford University, Stanford, CA, USA. She has led and contributed EU funded projects such as H2020 NRG5, eWINE, WISHFUL, FP7 CREW, Planetdata, ACTIVE, under various positions. She has advised/coadvised more than six M.Sc. and Ph.D. students. She has consulted public and private institutions. She has coauthored over 100 papers including in IEEE COMST, IEEE WICOM Magazine, IEEE OJCOMS and Access. Her research interest includes developing the next generation smart infrastructures that surround us and improve the quality of our life. Dr. Fortuna contributed to community work as a TPC member, the track chair, and a reviewer at several IEEE conferences including Globecom and ICC.

### v2

@@ Acknowledgments
This work was funded by the Slovenian Research and Innovation Agency (grant P2-0016) and by the European Union’s Horizon Europe Framework Program SNS-JU (grant agreement No. 101096456, NANCY). Instead of traditional spell-checking and language correction tools, some manually written paragraphs in the introduction and related work were refined using AI to improve flow and language quality.
@@
[Uncaptioned image] Blaž Bertalanič received his Ph.D. degree with highest distinction at the Faculty of Electrical engineering, University of Ljubljana. He is currently working as a researcher at Sensorlab, Jožef Stefan Institute. His main research interests are connected to advancement of machine learning and AI algorithms, especially in the context of time series analysis and smart infrastructures. Blaz is an IEEE member and with several leadership positions in the Slovenian chapter with over 15 IEEE publications.
[Uncaptioned image] George Floros is an Assistant Professor in the Department of Electronic & Electrical Engineering at Trinity College Dublin. Previously, he was a Postdoctoral Researcher and Lecturer at the Department of Electrical and Computer Engineering at the University of Thessaly, Greece. He received his Diploma in Engineering, MSc, and PhD degrees from the same department in 2013, 2015, and 2019, respectively. His research interests primarily focus on the fields of electronic design automation (EDA), semiconductor device modeling, circuit simulation, model order reduction, thermal analysis of ICs, as well as VLSI design techniques and embedded systems. Throughout his academic career, he has authored over 10 journal articles, 30 conference papers, 3 posters, and abstracts.
@@
[Uncaptioned image] Carolina Fortuna is a research associate professor with the Jožef Stefan Institute where she leads Sensorlab. She was a Post-Doctoral Researcher with Ghent University, Ghent, Belgium. She was a Visiting researcher in Infolab with Stanford University, Stanford, CA, USA. She has led and contributed EU funded projects such as H2020 NRG5, eWINE, WISHFUL, FP7 CREW, Planetdata, ACTIVE, under various positions. She has advised/coadvised more than six M.Sc. and Ph.D. students. She has consulted public and private institutions. She has coauthored over 100 papers including in IEEE COMST, IEEE WICOM Magazine, IEEE OJCOMS and Access. Her research interest includes developing the next generation smart infrastructures that surround us and improve the quality of our life. Dr. Fortuna contributed to community work as a TPC member, the track chair, and a TPC member at several IEEE conferences including Globecom and ICC.